If spray sprinklers are going to be used to protect a Class I commodity at a density of 0.15 gpm/...

  1. If spray sprinklers are going to be used to protect a Class I commodity at a density of 0.15 gpm/sq. ft what is the smallest orifice sprinkler that is permitted to be used?

    A.

    k-4.2

    B.

    k-5.6

    C.

    k-8.0

    D.

    k-11.2

    E.

    k-14.0
Answer #1

Generally, the flow rate of a spray sprinkler can be given by

q=k\\sqrt{p}

where k is the K-factor used in fire protection engineering,

p is the pressure at the sprinklers.

here we can have the smallest flow rate for finding smallest orifice sprinkler,thus we can have the calculation as follows

(a)k-4.2

q=4.2\\sqrt{.15} =10.844

(b)k=5.6

q=5.6\\sqrt{.15} =14.45

(c)k=8.0

q=8.0\\sqrt{.15} =20.665

(d)k=11.2

q=11.2\\sqrt{.15} =28.92

(e)k=14.0

q=14.0\\sqrt{.15} =36.14

Thus we can have the smallest flow rate 10.664 GPM at A.k=4.2 which can have smallest orifice sprinkler.